Status Report

SOHO Daily Meeting Minutes for Thursday, 20 June 2002

By SpaceRef Editor
June 20, 2002
Filed under , ,

* RHESSI target for today is AR 0001; target for tomorrow
is TBD.

* PDU Maintenance today at 12:40pm to 3:20pm local time.

* No NRT for up to an hour at 11am today for CELIAS patches.

* Voice Distribution System (VDS) will be down from 4:00 pm
June 22 (Saturday) until 12:00 am June 23 to install a power
supply subsystem.

* Commanding schedule for next week:

Monday – Gap until 14 UT then D27 all day.
Tuesday – Gap from 17-19 UT but good stations for the day.
Wedesday – D27 until 19 UT then D16.
Thursday – D27 all day.
Friday – Gap until 15 UT then D16 all day.
Saturday – Commanding is good all day.
Sunday – D16 until 16:30 UT then D27.

Spacecraft Status: Nominal

Spacecraft Anomalies: None

Accomplished Activities: VIRGO, CELIAS

Planned Activities: VIRGO, CELIAS, Nominal HGA Table,
CELIAS DPU Patch, Kevlar Cutter Switch
On (to reduce trickle charge on BATT2)

Upcoming Operations: None

Ground Anomalies:
170/1454 Lost 4.5 minutes of TLM due to a SCP failure at
D27. Data from 14:54:37 to 14:55:35 is
unrecoverable, rest can be recovered from the CDR.
171/0415 Invalid clock correlation output due to a TGC
failure at D66.

06/19 DOY 170 – 06/20 DOY 171

EIT daily On-Line Solar Status and observations:

LASCO daily On-Line Solar Status and observations:

Web Page for Planning:

CDS: Nominal. Loop studies on AR 9992 today. Friday –
CDS will return to the AR 0001 pointing.

UVCS: Nominal. Cassini conjunction and mini-synoptic for today
and tomorrow.

LASCO: Nominal. C2/C3 synoptics and JOP159. The stray light
pattern in C3 will be subtracted out within a week.

EIT: Nominal. 195 CME watch and synoptics.

MDI: Nominal. High resolution dopplergrams and magnetograms
for continuous contact.

TRACE: Nominal. Continue observing AR0001.

SpaceRef staff editor.